OpenStudy (anonymous):

factor completely 16c^2-121

OpenStudy (anonymous):

Use \[ A^2 - B^2 =(A-B)(A+B) \]

OpenStudy (anonymous):

\[ 16c^2-121 =(4 c -11)(4c+11) \]

OpenStudy (anonymous):